### v1

# Annealing Approximation in Master-Node Network Model

Abstract
This paper investigates absorbing-state phase transitions in opinion dynamics through a master-node network model analyzed using annealing approximation. We develop a theoretical framework examining three fundamental regimes: systems converging to complete disagreement, complete consensus, or both states depending on initial conditions. The phase behavior is governed by two key chiral parameters - R measuring right-oriented influence and L measuring left-oriented influence in the network interactions. Our analysis reveals a rich phase diagram featuring both continuous and discontinuous transitions between disordered and ordered phases. The discontinuous transition emerges in systems with two absorbing states, where the final configuration depends critically on initial opinion distributions. The annealing approximation provides fundamental insights into how asymmetric social influences (chirality) shape collective opinion formation, acting as a symmetry-breaking element that drives the system toward polarization or consensus.
1 Introduction
The study of absorbing-state transitions provides fundamental insights into opinion dynamics within social systems. These transitions occur when a system evolves into a stable configuration from which it cannot escape, representing either consensus formation or the dominance of a particular opinion among agents. This phenomenon is particularly relevant for understanding how opinions emerge, propagate, and stabilize in populations, revealing the mechanisms through which individuals influence each other and converge toward shared beliefs. In social contexts, absorbing states represent conditions where opinions become fixed, with agents unable to return to previous belief states. This process is crucial for understanding consensus formation, where groups reach stable agreements shaped by social interactions, cultural influences, and external pressures [1, 2, 3, 4].
Models of opinion formation demonstrate how individuals modify their beliefs through peer interactions, leading to either consensus or polarization [5, 6]. A critical aspect of these dynamics involves the interplay between individual cognitive processes and broader sociocultural contexts, which provide frameworks for interpreting and responding to social cues [6, 7]. Absorbing-state phase transitions, where systems become trapped in particular configurations, offer valuable perspectives on how individuals mutually influence each other and converge toward collective beliefs [8, 9]. The critical properties observed in these transitions suggest universal behavior across diverse network structures, with local interactions playing a pivotal role in determining outcomes [10].
Cultural factors significantly shape these dynamics by embedding norms and values into interaction processes, thereby influencing the trajectory of opinion evolution [11, 12]. Understanding these mechanisms is essential for explaining how stable opinion patterns emerge in societies, revealing the complex relationships between individual beliefs, social influence, and cultural context [13, 14]. Such insights extend beyond technical modeling, providing a comprehensive framework for analyzing how social systems stabilize and evolve.
Recent developments in opinion dynamics have introduced master-node models to investigate multi-agent systems influenced by central nodes within networked structures [15, 16]. The work of [15] examines how a master node with preferential opinions affects consensus formation and social dynamics. Using stochastic quenched disorder models and mean-field approximations, the study reveals a rich phase diagram connecting critical parameters with collective behavior transitions, offering insights into mechanisms underlying morality, innovation, and cooperation in social systems. Complementing this, [16] conducts rigorous analysis of critical exponents in one-dimensional networks through Monte Carlo simulations and finite-size scaling methods. The research identifies a continuous phase transition to an absorbing state that does not conform to known universality classes, highlighting the model’s unique properties.
These studies provide complementary perspectives on networked agent dynamics under master node influence. While [15] establishes the foundational framework and demonstrates phase transitions as key mechanisms in collective behavior, [16] offers detailed characterization of critical properties through exponent analysis. Together, they advance understanding of decentralized decision-making processes while contributing to non-equilibrium statistical mechanics. The current work builds upon these foundations by introducing a chiral master node formulation that incorporates geometric properties and Berry phase effects, enabling more comprehensive analysis of symmetry breaking patterns in opinion dynamics. The general formulation of the model, as shown in Section 2, contains so many parameters that it leads to a system that is not analytically accessible at first sight. Nevertheless, introducing an annealing approximation (on the variable associated with the influence of the master node) and examining the model at the mean-field level generates a system where several relevant cases can be characterized by just two quantities, as shown in Section 3. The analysis is then organized into three parts (Sections 4 to 6) and the role of the master node in this approximation is discussed in Section 7. Some final remarks can be found in Section 8.
2 General formulation (continuous approach)
The state of the above-defined model is completely described by the vector $s:=\begin{pmatrix}s_{1},&\cdots,&s_{N}\end{pmatrix}$, where $s_{i}\in\{0,1\}$ for each one of the $N$ vertices of the model. The dynamics is governed by the master equation
$\frac{\textup{d}}{\textup{d}t}P(s,t)=\sum_{i=1}^{N}\left[w_{i}(s^{i})P(s^{i},t)-w_{i}(s)P(s,t)\right],$ (1)
which characterizes the time evolution of the probability $P$ of the system begin at state $s$ at time $t$. In (1), $s^{i}:=\begin{pmatrix}s_{1},&\cdots,&1-s_{i},&\cdots,&s_{N}\end{pmatrix}$ stands for the state where the $i$-th state is replaced by its complement and $w_{i}$ is the transition rate (there should be no confusion with the parameters $w_{1,2,3,4}$ introduced in the model) of the individual at site $i$.
Neighbors $s_{i}\rightarrow s_{i}^{\prime}$  ($\Gamma_{i}\in\{0,1\}$)
$s_{i-1}$ $s_{i+1}$ $0\rightarrow 0$ $0\rightarrow 1$ $1\rightarrow 0$ $1\rightarrow 1$
00 00 $p_{1}+\Gamma_{i}\left(q_{1}-p_{1}\right)$ $\overline{p_{1}}+\Gamma_{i}\left(\overline{q_{1}}-\overline{p_{1}}\right)$ $p_{2}+\Gamma_{i}\left(q_{2}-p_{2}\right)$ $\overline{p_{2}}+\Gamma_{i}\left(\overline{q_{2}}-\overline{p_{2}}\right)$
$1$ $1$ $p_{3}+\Gamma_{i}\left(q_{3}-p_{3}\right)$ $\overline{p_{3}}+\Gamma_{i}\left(\overline{q_{3}}-\overline{p_{3}}\right)$ $p_{4}+\Gamma_{i}\left(q_{4}-p_{4}\right)$ $\overline{p_{4}}+\Gamma_{i}\left(\overline{q_{4}}-\overline{p_{4}}\right)$
00 $1$ $w_{1}+\Gamma_{i}\left(v_{1}-w_{1}\right)$ $\overline{w_{1}}+\Gamma_{i}\left(\overline{v_{1}}-\overline{w_{1}}\right)$ $w_{2}+\Gamma_{i}\left(v_{2}-w_{2}\right)$ $\overline{w_{2}}+\Gamma_{i}\left(\overline{v_{2}}-\overline{w_{2}}\right)$
$1$ 00 $w_{3}+\Gamma_{i}\left(v_{3}-w_{3}\right)$ $\overline{w_{3}}+\Gamma_{i}\left(\overline{v_{3}}-\overline{w_{3}}\right)$ $w_{4}+\Gamma_{i}\left(v_{4}-w_{4}\right)$ $\overline{w_{4}}+\Gamma_{i}\left(\overline{v_{4}}-\overline{w_{4}}\right)$
Table 1: Transition rates for all possible neighbor configurations $(s_{i-1},s_{i+1})$ and state changes $(s_{i}\rightarrow s_{i}^{\prime})$. The rates depend on both the base parameters ($p_{j}$, $w_{j}$, $v_{j}$, $q_{j}$) and the master node influence $\Gamma_{i}$. The overline notation $\overline{x}$ denotes the complementary probability $1-x$.
Following the rules presented in table 1, the transition rate can be represented by
$w_{i}(s)$ $=s_{i}\big{\{}\left(1-s_{i-1}\right)\left(1-s_{i+1}\right)\left[p_{2}+\Gamma_{i}\left(q_{2}-p_{2}\right)\right]+s_{i-1}s_{i+1}\left[p_{4}+\Gamma_{i}\left(q_{4}-p_{4}\right)\right]+$
$+\left(1-s_{i-1}\right)s_{i+1}\left[w_{2}+\Gamma_{i}\left(v_{2}-w_{2}\right)\right]+s_{i-1}\left(1-s_{i+1}\right)\left[w_{4}+\Gamma_{i}\left(v_{4}-w_{4}\right)\right]\big{\}}+$
$+\left(1-s_{i}\right)\big{\{}\left(1-s_{i-1}\right)\left(1-s_{i+1}\right)\left[\overline{p_{1}}+\Gamma_{i}\left(\overline{q_{1}}-\overline{p_{1}}\right)\right]+s_{i-1}s_{i+1}\left[\overline{p_{3}}+\Gamma_{i}\left(\overline{q_{3}}-\overline{p_{3}}\right)\right]+$
$+\left(1-s_{i-1}\right)s_{i+1}\left[\overline{w_{1}}+\Gamma_{i}\left(\overline{v_{1}}-\overline{w_{1}}\right)\right]+s_{i-1}\left(1-s_{i+1}\right)\left[\overline{w_{3}}+\Gamma_{i}\left(\overline{v_{3}}-\overline{w_{3}}\right)\right]\big{\}}$
$=s_{i}\left[A_{i}^{(2)}+\left(B_{i}^{(4)}-A_{i}^{(2)}\right)s_{i-1}+\left(B_{i}^{(2)}-A_{i}^{(2)}\right)s_{i+1}+\left(A_{i}^{(2)}+A_{i}^{(4)}-B_{i}^{(2)}-B_{i}^{(4)}\right)s_{i-1}s_{i+1}\right]+$
$+\left(1-s_{i}\right)\left[\overline{A_{i}^{(1)}}+\left(\overline{B_{i}^{(3)}}-\overline{A_{i}^{(1)}}\right)s_{i-1}+\left(\overline{B_{i}^{(1)}}-\overline{A_{i}^{(1)}}\right)s_{i+1}+\left(\overline{A_{i}^{(1)}}+\overline{A_{i}^{(3)}}-\overline{B_{i}^{(1)}}-\overline{B_{i}^{(3)}}\right)s_{i-1}s_{i+1}\right],$ (2)
where
$A_{i}^{(j)}:=p_{j}+\Gamma_{i}\left(q_{j}-p_{j}\right),\quad\overline{A_{i}^{(j)}}:=\overline{p_{j}}+\Gamma_{i}\left(\overline{q_{j}}-\overline{p_{j}}\right),\quad B_{i}^{(j)}:=w_{j}+\Gamma_{i}\left(v_{j}-w_{j}\right)\;\text{ and }\;\overline{B_{i}^{(j)}}:=\overline{w_{j}}+\Gamma_{i}\left(\overline{v_{j}}-\overline{w_{j}}\right)$ (3)
for $j\in\{1,2,3,4\}$.
The magnetization of the $i$-th individual is defined as usual,
$m_{i}(t):=\sum_{s}s_{i}P(s,t),$ (4)
where the sum runs over the set $\{0,1\}^{N}$. From the master equation (1), the dynamics of the magnetization is described by
$\frac{\textup{d}}{\textup{d}t}m_{i}(t):=\sum_{s}\left(1-2s_{i}\right)w_{i}(s)P(s,t),$ (5)
and from the particular form (2), it implies
$\frac{\textup{d}}{\textup{d}t}m_{i}(t)$ $=\overline{A_{i}^{(1)}}-\left(\overline{A_{i}^{(1)}}+A_{i}^{(2)}\right)m_{i}(t)+\left(\overline{B_{i}^{(3)}}-\overline{A_{i}^{(1)}}\right)m_{i-1}(t)+\left(\overline{B_{i}^{(1)}}-\overline{A_{i}^{(1)}}\right)m_{i+1}(t)+$
$+\left(\overline{A_{i}^{(1)}}+A_{i}^{(2)}-\overline{B_{i}^{(3)}}-B_{i}^{(4)}\right)q_{i,i-1}(t)+\left(\overline{A_{i}^{(1)}}+A_{i}^{(2)}-\overline{B_{i}^{(1)}}-B_{i}^{(2)}\right)q_{i,i+1}(t)+$
$+\left(\overline{A_{i}^{(1)}}+\overline{A_{i}^{(3)}}-\overline{B_{i}^{(1)}}-\overline{B_{i}^{(3)}}\right)q_{i-1,i+1}(t)+$
$+\left(\overline{B_{i}^{(1)}}+B_{i}^{(2)}+\overline{B_{i}^{(3)}}+B_{i}^{(4)}-\overline{A_{i}^{(1)}}-A_{i}^{(2)}-\overline{A_{i}^{(3)}}-A_{i}^{(4)}\right)c_{i-1,i,i+1}(t),$ (6)
where
$q_{i,j}(t):=\sum_{s}s_{i}s_{j}P(s,t)\;\textnormal{ and }\;c_{i,j,k}(t):=\sum_{s}s_{i}s_{j}s_{k}P(s,t).$ (7)
3 Annealed mean-field analysis
The usual assumptions
$m_{i}(t)=m(t),\quad q_{i,j}(t)=m^{2}(t)\quad(i\neq j),\quad\text{ and }\quad c_{i,j,k}(t)=m^{3}(t)\quad(i\neq j\neq k\neq i),$ (8)
are introduced to investigate the dynamics (6) at the mean-field level. Furthermore, the annealed approach is assumed for the random variable $\Gamma_{i}$, which is then replaced by its expected value $\Gamma_{i}\rightarrow\langle\Gamma_{i}\rangle=\rho$, which is the fraction of connections from the master node. As a consequence of these hypothesis, all the subindices of $A_{i}^{(j)}\rightarrow A^{(j)}$, $\overline{A_{i}^{(j)}}\rightarrow\overline{A^{(j)}}$, $B_{i}^{(j)}\rightarrow B^{(j)}$ and $\overline{B_{i}^{(j)}}\rightarrow\overline{B^{(j)}}$ are now irrelevant and are going to be omitted. Furthermore, the dependence of $m(t)$ on time is also suppressed from now on in the notation.
Taking into account these approximations, it is straightforward to show that the dynamics can now be described by
$\dot{m}=F(m),$ (9)
where $\dot{m}$ stands for the time derivative of $m$ and
$F(m):=\overline{A^{(1)}}-\left(3\overline{A^{(1)}}+L\right)m+\left(3\overline{A^{(1)}}+R+2L\right)m^{2}-\left(R+L+\overline{A^{(1)}}+A^{(4)}\right)m^{3}.$ (10)
Here, the quantities $L$ and $R$ are defined as
$L:=A^{(2)}-\overline{B^{(1)}}-\overline{B^{(3)}}\quad\text{ and }\quad R:=\overline{A^{(3)}}-B^{(2)}-B^{(4)}$ (11)
and play a major role in the analysis of the system.
At this point, it is interesting to restrict the analysis where some well-knowm absorbing states are present in the system, namely the stationary states $m^{\ast}=0$ and/or $m^{\ast}=1$ in the solution of (9). From (10), one can see that they correspond to the assumption $\overline{A^{(1)}}=0$ and $A^{(4)}=0$, respectively. In what follows, the analysis is divided into the following cases:
Case I $m^{\ast}=0$ is an absorbing state but $m^{\ast}=1$ is not
Case II $m^{\ast}=1$ is an absorbing state but $m^{\ast}=0$ is not
Case III $m^{\ast}=0$ and $m^{\ast}=1$ are both absorbing states
4 System with absorbing state $m^{\ast}=0$ (without the absorbing state $m^{\ast}=1$) – Case I
In this scenario, one has $\overline{A^{(1)}}=0$, and the equation of motion is cast as $\dot{m}=F_{0}(m)$, where
$F_{0}(m)$ $:=-Lm+\left(R+2L\right)m^{2}-\left(R+L+A^{(4)}\right)m^{3}$
$=-m\left[L-\left(R+2L\right)m+\left(R+L+A^{(4)}\right)m^{2}\right].$ (12)
As seen from (12), apart from the stationary point $m^{\ast}=0$, two other stationary solutions, $m_{+}^{(0)}$ and $m_{-}^{(0)}$ , can emerge depending on the choice of the pair $(L,R)$. These values are given by
$m_{\pm}^{(0)}:=\frac{R+2L\pm\sqrt{R^{2}-4LA^{(4)}}}{2\left(R+L+A^{(4)}\right)}.$ (13)
The stationary points in this case are
$m^{\ast}\in\{0,m_{\pm}^{(0)}\}.$ (14)
If $A^{(4)}=0$, the system will also have $m^{\ast}=1$ as one of its stationary points. Since this case (00 and $1$ being both fixed points of the dynamical system) will be examined later, it will be assumed that $A^{(4)}\neq 0$ (or $A^{(4)}>0$) in this section. The stability analysis is going to be performed through the (one-dimensional) isocline method, which stems on the analysis of the sign of (12). Consequently, the relative positions of the points 00, $m_{+}^{(0)}$ and $m_{-}^{(0)}$ plays a major rule in this procedure and is detailed in the Appendix B, and the results are summarized in figure 1.
Figure 1: Diagram $R\times L$ for case I, where $m^{\ast}=0$ is an absorbing state, while $m=1$ is not.
Defining the regions
$\left\{\begin{array}[]{ll}\Omega_{0}^{(0)}&:=\left\{(L,R)\in\mathbb{R}^{2}:L\geq 0\text{ and }R<\sqrt{4LA^{(4)}}\right\}\cup\left\{(0,0)\right\}\\ \Omega_{+}^{(0)}&:=\left\{(L,R)\in\mathbb{R}^{2}:L<0\right\}\cup\left\{(L,R)\in\mathbb{R}^{2}:L=0\text{ and }R>0\right\}\\ \Omega_{0+}^{(0)}&:=\left\{(L,R)\in\mathbb{R}^{2}:L>0\text{ and }R\geq\sqrt{4LA^{(4)}}\right\}\end{array}\right.,$ (18)
with $\Omega_{0}^{(0)}\cup\Omega_{+}^{(0)}\cup\Omega_{0+}^{(0)}=\mathbb{R}^{2}$. As shown in figure 1, the regions $\Omega_{0}^{(0)}$ and $\Omega_{+}^{(0)}$ contain a single stationary point, which are $m^{\ast}=0$ and $m^{\ast}=m_{+}^{(0)}$, respectively. On the other hand, the region $\Omega_{0+}^{(0)}$ displays two stationary points, 00 and $m_{+}^{(0)}$; the initial condition $m(t=0)=m_{0}$ determines the trajectory toward one of these two fixed points in $\Omega_{0+}^{(0)}$. In Appendix B, it is shown that $10<m_{-}^{(0)}<m_{+}^{(0)}<$ 1 in this region, and
$\lim_{t\rightarrow\infty}m(t)=\left\{\begin{array}[]{lcl}0&,&m_{0}<m_{-}^{(0)}\\ m_{+}^{(0)}&,&m_{0}>m_{-}^{(0)}\end{array}\right..$ (21)
In other words, the point $m=m_{-}^{(0)}$, which is an unstable fixed point, is a separatrix between these two stationary points, 00 and $m_{+}^{(0)}$.
Although the region $\{(L,R):\mathbb{R}^{2}:L=0\text{ and }R<0\}$ belongs to $\Omega_{0}^{(0)}$, the transition to the region $\Omega_{+}^{(0)}$ is smooth in the sense that
$\lim_{\underset{R\leq 0}{L\rightarrow 0^{-}}}m_{+}^{(0)}=0$ (22)
from (13). Conversely,
$\lim_{\underset{R>0}{L\rightarrow 0^{-}}}m_{+}^{(0)}=\frac{R}{R+A^{(4)}}>0$ (23)
in the boundary between $\Omega_{+}^{(0)}$ and $\Omega_{0+}^{(0)}$.
In the frontier between $\Omega_{0}^{(0)}$ and $\Omega_{0+}^{(0)}$, one has
$\lim_{\underset{R>0}{L\rightarrow\left(\nicefrac{{R^{2}}}{{4A^{(4)}}}\right)^{-}}}m_{+}^{(0)}=\frac{R}{R+2A^{(4)}}>0,$ (24)
indicating a discontinuity of $m_{+}^{(0)}$ when crossing this boundary.
Summarizing, if one regards $m_{+}^{(0)}$ as an order parameter, the transition is of first-order between $\Omega_{0}^{(0)}$ and $\Omega_{0+}^{(0)}$, continuous between $\Omega_{+}^{(0)}$ and $\Omega_{0}^{(0)}$ or $\Omega_{0+}^{(0)}$. The origin $(L,R)=(0,0)$ is then a tricritical point.
5 System with absorbing state $m^{\ast}=1$ (without the absorbing state $m^{\ast}=0$) – Case II
In this scenario, one has $A^{(4)}=0$ and $\overline{A^{(1)}}=0$, and the equation of motion is cast as $\dot{m}=F_{1}(m)$, where
$F_{1}(m)$ $:=\overline{A^{(1)}}-\left(3\overline{A^{(1)}}+L\right)m+\left(3\overline{A^{(1)}}+R+2L\right)m^{2}-\left(R+L+\overline{A^{(1)}}\right)m^{3}$
$=\left(1-m\right)\left[\overline{A^{(1)}}-\left(2\overline{A^{(1)}}+L\right)m+\left(R+L+\overline{A^{(1)}}\right)m^{2}\right]$ (25)
Two more stationary points, $m_{+}^{(1)}$ and $m_{-}^{(1)}$, are potentially available, depending on the choice of the parameters, and are given by
$m_{\pm}^{(1)}:=\frac{2\overline{A^{(1)}}+L\pm\sqrt{L^{2}-4R\overline{A^{(1)}}}}{2\left(R+L+\overline{A^{(1)}}\right)}.$ (26)
The stationary points in this case are
$m^{\ast}\in\{1,m_{\pm}^{(1)}\}.$ (27)
Analogously in the previous section, the case $\overline{A^{(1)}}=0$ is excluded in the analysis here, since it will be explored in the next section. Moreover, the technical details of the stationary analysis of 25 is shown in Appendix C, and the results can be summarized in the diagram of figure 2.
Figure 2: Diagram $R\times L$ for case II, where $m^{\ast}=1$ is an absorbing state, while $m=0$ is not.
Defining the regions
$\left\{\begin{array}[]{ll}\Omega_{1}^{(1)}&:=\left\{(L,R)\in\mathbb{R}^{2}:R\geq 0\text{ and }L<\sqrt{4R\overline{A^{(1)}}}\right\}\cup\left\{(0,0)\right\}\\ \Omega_{-}^{(1)}&:=\left\{(L,R)\in\mathbb{R}^{2}:R<0\right\}\cup\left\{(L,R)\in\mathbb{R}^{2}:R=0\text{ and }L>0\right\}\\ \Omega_{1-}^{(1)}&:=\left\{(L,R)\in\mathbb{R}^{2}:R>0\text{ and }L\geq\sqrt{4R\overline{A^{(1)}}}\right\}\end{array}\right.,$ (31)
with $\Omega_{1}^{(1)}\cup\Omega_{-}^{(1)}\cup\Omega_{1-}^{(1)}=\mathbb{R}^{2}$. As shown in figure 2, the regions $\Omega_{1}^{(1)}$ and $\Omega_{-}^{(1)}$ contain a single stationary point, which are $m^{\ast}=1$ and $m^{\ast}=m_{-}^{(1)}$, respectively. On the other hand, the region $\Omega_{1-}^{(1)}$ displays two stationary points, $1$ and $m_{-}^{(1)}$; the initial condition $m(t=0)=m_{0}$ determines the trajectory toward one of these two fixed points in $\Omega_{1-}^{(1)}$. In Appendix C, it is shown that $10<m_{-}^{(1)}<m_{+}^{(1)}<$ 1 in this region, and
$\lim_{t\rightarrow\infty}m(t)=\left\{\begin{array}[]{lcl}1&,&m_{0}>m_{+}^{(1)}\\ m_{-}^{(0)}&,&m_{0}<m_{+}^{(1)}\end{array}\right..$ (34)
In other words, the point $m=m_{+}^{(1)}$, which is an unstable fixed point, is a separatrix between these two stationary points, $1$ and $m_{-}^{(1)}$.
Although the region $\{(L,R):\mathbb{R}^{2}:R=0\text{ and }L<0\}$ belongs to $\Omega_{1}^{(1)}$, the transition to the region $\Omega_{-}^{(1)}$ is smooth in the sense that
$\lim_{\underset{L\leq 0}{R\rightarrow 0^{-}}}m_{-}^{(1)}=1$ (35)
from (26). Conversely,
$\lim_{\underset{L>0}{R\rightarrow 0^{-}}}m_{-}^{(1)}=\frac{\overline{A^{(1)}}}{R+\overline{A^{(1)}}}>0$ (36)
in the boundary between $\Omega_{-}^{(1)}$ and $\Omega_{1-}^{(1)}$.
In the frontier between $\Omega_{1}^{(1)}$ and $\Omega_{1-}^{(1)}$, one has
$\lim_{\underset{L>0}{R\rightarrow\left(\nicefrac{{L^{2}}}{{4\overline{A^{(1)}}}}\right)^{-}}}m_{-}^{(1)}=\frac{2\overline{A^{(1)}}}{2\overline{A^{(1)}}+L}>0,$ (37)
indicating a discontinuity of $m_{-}^{(1)}$ when crossing this boundary.
Summarizing, if one regards $m_{-}^{(1)}$ as an order parameter, the transition is of first-order between $\Omega_{1}^{(1)}$ and $\Omega_{1-}^{(1)}$, continuous between $\Omega_{-}^{(1)}$ and $\Omega_{1}^{(1)}$ or $\Omega_{1-}^{(1)}$. The origin $(L,R)=(0,0)$ is then a tricritical point.
6 System with absorbing state $m^{\ast}=0$ and $m^{\ast}=1$ – Case III
In this last scenario, one has $A^{(4)}=0$ and $\overline{A^{(1)}}=0$, and the equation of motion is cast as $\dot{m}=F_{01}(m)$, where
$F_{01}(m)$ $:=-Lm+\left(R+2L\right)m^{2}-\left(R+L\right)m^{3}$
$=m\left(1-m\right)\left[\left(R+L\right)m-L\right],$ (38)
and the stationary points $m^{\ast}=0$ and $m^{\ast}=1$ are naturally present. Apart from them, it is also possible to have a third one,
$\tilde{m}^{(01)}:=\frac{L}{R+L},$ (39)
depending on the choice of the pair $(L,R)$. The set of stationary points in this case are, therefore,
$m^{\ast}\in\{0,1,\tilde{m}^{(01)}\}.$ (40)
The stability analysis is summarized in figure 3, while the technical details are found in Appendix D.
Figure 3: Diagram $R\times L$ for case III, where $m^{\ast}=0$ and $m^{\ast}=1$ are absorbing states.
Defining the regions
$\left\{\begin{array}[]{ll}\Omega_{0}^{(01)}&:=\left\{(L,R)\in\mathbb{R}^{2}:L\geq 0\text{ and }R\leq 0\right\}\setminus\{(0,0)\}\\ \Omega_{1}^{(01)}&:=\left\{(L,R)\in\mathbb{R}^{2}:L\leq 0\text{ and }R\geq 0\right\}\setminus\{(0,0)\}\\ \Omega_{\tilde{m}}^{(01)}&:=\left\{(L,R)\in\mathbb{R}^{2}:L<0\text{ and }R<0\right\}\\ \Omega_{01}^{(01)}&:=\left\{(L,R)\in\mathbb{R}^{2}:L>0\text{ and }R>0\right\}\end{array}\right.,$ (45)
with $\Omega_{0}^{(01)}\cup\Omega_{1}^{(01)}\cup\Omega_{\tilde{m}}^{(01)}\cup\Omega_{01}^{(01)}=\mathbb{R}^{2}\setminus\{(0,0)\}$. As shown in figure 3, the regions $\Omega_{0}^{(01)}$, $\Omega_{1}^{(01)}$ and $\Omega_{\tilde{m}}^{(01)}$ contain a single stationary point, which are $m^{\ast}=0$, $m^{\ast}=1$ and $m^{\ast}=\tilde{m}^{(01)}$, respectively. On the other hand, the region $\Omega_{01}^{(01)}$ displays two stationary points, 00 and $1$; the initial condition $m(t=0)=m_{0}$ determines the trajectory toward one of these two fixed points in $\Omega_{01}^{(01)}$. In Appendix D, it is shown that $10<\tilde{m}^{(01)}<$ 1 in this region, and
$\lim_{t\rightarrow\infty}m(t)=\left\{\begin{array}[]{lcl}0&,&m_{0}<\tilde{m}^{(01)}\\ 1&,&m_{0}>\tilde{m}^{(01)}\end{array}\right..$ (48)
In other words, the point $m=\tilde{m}^{(01)}$, which is an unstable fixed point, is a separatrix between these two stationary points, 00 and $1$.
Although the region $\{(L,R):\mathbb{R}^{2}:R=0\text{ and }L<0\}$ belongs to $\Omega_{1}^{(01)}$, and $\{(L,R):\mathbb{R}^{2}:L=0\text{ and }R<0\}$ belongs to $\Omega_{0}^{(01)}$, the transition between $\Omega_{\tilde{m}}^{(01)}$ and these involved regions are smooth in the sense that
$\lim_{\underset{L<0}{R\rightarrow 0^{-}}}\tilde{m}^{(01)}=1\quad\textnormal{ and }\quad\lim_{\underset{R<0}{L\rightarrow 0^{-}}}\tilde{m}^{(01)}=0.$ (49)
from (39).
The stationary behviour of the point $(0,0)$ should be examined separately, since the equation of motion is $\dot{m}=0$ in this point when $A^{(4)}=\overline{A^{(1)}}=0$. This means that no dynamics takes place at this particular point and the stationary value for $m(t)$ equals the initial value.
7 The role of the master node
There are some general observations concerning the variables
$L:=A^{(2)}-\overline{B^{(1)}}-\overline{B^{(3)}}\quad\textnormal{ and }\quad R:=\overline{A^{(3)}}-B^{(2)}-B^{(4)},$ (50)
where the involved quantities (defined before) are stated here for clarity in table 2.
$010$ $\overset{A^{(2)}}{\longrightarrow}$ $000$ $101$ $\overset{\overline{A^{(3)}}}{\longrightarrow}$ $111$
$001$ $\overset{\overline{B^{(1)}}}{\longrightarrow}$ $011$ $011$ $\overset{B^{(2)}}{\longrightarrow}$ $001$
$100$ $\overset{\overline{B^{(3)}}}{\longrightarrow}$ $110$ $110$ $\overset{B^{(4)}}{\longrightarrow}$ $100$
Table 2: Transition rates $A^{(2)}$, $\overline{B^{1}}$, $\overline{B^{3}}$, $\overline{A^{3}}$, $B^{(2)}$ and $B^{(4)}$.
Some general observations can then be made from these transition rates.
  1. 1.
    Tendency to state $m=0$: $L>0$ and $R<0$, as a consequence of large $A^{(2)}$, $B^{(2)}$ and $B^{(4)}$, and low $\overline{A^{(3)}}$, $\overline{B^{(1)}}$ and $\overline{B^{(3)}}$. This is valid even $m=0$ not being an absorbing state.
  2. 2.
    Tendency to state $m=1$: $L<0$ and $R>0$, as a consequence of low $A^{(2)}$, $B^{(2)}$ and $B^{(4)}$, and large $\overline{A^{(3)}}$, $\overline{B^{(1)}}$ and $\overline{B^{(3)}}$. This is valid even $m=1$ not being an absorbing state.
  3. 3.
    Tendency to cluster formation: $L>0$ and $R>0$, as a consequence of large $A^{(2)}$ and $\overline{A^{(3)}}$.
  4. 4.
    Tendency to cluster destruction: $L<0$ and $R<0$, as a consequence of low $A^{(2)}$ and $\overline{A^{(3)}}$.
Note that these four cases are not disjoint, but one can see the potential impact of the master node in the dynamics. To facilitate the analysis, let us consider two extremal cases. Firstly, if the master node is decoupled by letting $\Gamma_{i}=0$ for all sites in (3), it leads to
$L(\Gamma_{i}=0)=p_{2}-\overline{w_{1}}-\overline{w_{3}}\quad\textnormal{ and }\quad R(\Gamma_{i}=0)=\overline{p_{3}}-w_{2}-w_{4}.$ (51)
On the other hand, the other extremal case, $\Gamma_{i}=1$ for all vertices $i$, leads to
$L(\Gamma_{i}=1)=q_{2}-\overline{v_{1}}-\overline{v_{3}}\quad\textnormal{ and }\quad R(\Gamma_{i}=1)=\overline{q_{3}}-v_{2}-v_{4}.$ (52)
From (51) and (52), all the involved transition rates are independent. This means that starting from a setup where the master node has no role ($\Gamma_{i}=0$), one can move to another scenario by letting the master node participate in the dynamics. As an example, one might choose the parameters such that the dynamics tends to the state $m=0$, but turning on the master node leads the system to the state $m=1$.
8 Discussion
Model Analysis and Key Findings
The master-node network model aims to provide a comprehensive framework for understanding opinion dynamics. The model can be used as a general framework in the spirit of other already proposed frameworks [17, 18, 19]. Here, we saw that mean field analysis show that a significant part of the parameter space can be understood through three distinct regimes of absorbing states. Indeed, by choosing the parameters properly, the model can represent several traditional effects from opinion mechanics, like contrarians or inflexible. Here, by employing annealing approximation and mean-field analysis, we have reduced the complex many-body problem to a tractable form while preserving essential features of the dynamics. The most significant theoretical advance lies in the complete characterization of phase diagrams for all three cases - systems converging to complete disagreement ($m^{*}=0$), complete consensus ($m^{*}=1$), or both states depending on initial conditions. The identification of tricritical points at $(L,R)=(0,0)$ reveals fundamental limitations in our ability to connect disordered and ordered phases smoothly.
Interpretation of R and L Parameters
The quantities $R$ and $L$ emerge as fundamental order parameters representing competing directional influences in the network. However, they do not correspond solely to movements towards one choice or another, as there are positive and negative terms in each of them. And yet, physically, a larger $R$ always corresponds to a right-oriented influence, combining the tendency to adopt right-neighbor opinions against resistance to such adoption. Conversely, $L$ quantifies left-oriented influence through similar mechanisms. Their ratio determines the system’s chiral character, with $R>L$ favoring right-biased consensus and vice versa. These parameters capture the inherent asymmetry in opinion propagation through the network. Interestingly, depending on those quantities, we observe a fixed point with coexistence of both opinions, like previous models with inflexibles [20, 21] or with asymmetric influences [galam2004dynamic].
Phase Transitions and Critical Behavior Let us analyze separately the scenarios with a single trivial stationary point (cases I and II) and the case where two trivial fixed points are displayed (case III). In the former, three regions, $\Omega_{trivial}$ (corresponding to $\Omega_{0}^{(0)}$ and $\Omega_{1}^{(1)}$ in cases I and II, respectively), $\Omega_{nontrivial}$ (domains $\Omega_{+}^{(0)}$ and $\Omega_{-}^{(1)}$ in cases I and II, in this order) and $\Omega_{2}$ (the regions $\Omega_{0+}^{(0)}$ and $\Omega_{1-}^{(1)}$ in cases I and II, respectively) can be identified. The transition from $\Omega_{trivial}$ to $\Omega_{nontrivial}$ is smooth in the sense of (22) and (35), while a discontinuity is observed in the frontier between $\Omega_{nontrivial}$ and $\Omega_{2}$. Therefore, the point $(0,0)$ can be regarded as a tricritical point. The region $\Omega_{2}$ has two fixed points, where the initial condition of the dynamics plays a major role in determining to which one the system converges; furthermore, the boundary that separates to which stationary point the systems moves toward to is given by an unstable stationary point (see (21) and (34)). Finally, the transition from $\Omega_{nontrivial}$ and $\Omega_{2}$ is smooth because it presents the behavior associated with the nontrivial stationary point, shared by these two regions. On the other hand, in case III, two trivial stationary points ($m^{\ast}=0$ and $m^{\ast}=1$) are present and generate the domains $\Omega_{0}^{(01)}$ and $\Omega_{1}^{(01)}$. This time, the boundaries of these two domains are closed (in the sense that the stationary value of $m(t)$ matches the value at its core), except for the origin $(0,0)$, where the stationary point equals the initial value. The region $\Omega_{01}^{(01)}$, which is the analogous phase of $\Omega_{2}$, displays a stationary point according to the initial condition.
Implications and Future Directions These results provide significant insights for modeling real-world social systems, particularly in understanding how centralized influences can alter collective behavior. The phase diagrams offer predictive power for consensus formation. Future work should investigate the model beyond mean-field approximation and explore connections to measurable social quantities in empirical networks. The framework also suggests principles for engineering consensus in organizational systems.
Conclusion
This work has established a comprehensive theoretical framework for analyzing absorbing-state transitions in opinion dynamics through a master-node network model. The annealing approximation approach has revealed several fundamental insights about collective behavior in social systems:
1) The model demonstrates how simple interaction rules can generate complex phase diagrams with both continuous and discontinuous transitions between disordered and ordered states. The characterization of three distinct regimes (single absorbing state at $m^{*}=0$, at $m^{*}=1$, and bistable systems) provides a complete taxonomy of possible consensus formation scenarios.
2) The chiral parameters $R$ and $L$ emerge as powerful descriptors of system behavior, quantifying the inherent asymmetry in opinion propagation through directional influences. Their ratio determines the ultimate consensus state and helps explain polarization phenomena.
3) The identification of tricritical points and the detailed analysis of transition lines offer practical tools for predicting when small changes in network parameters may lead to dramatic shifts in collective opinion.
4) The master node’s role as a symmetry-breaking element suggests design principles for influencing real-world social systems, from organizational management to public opinion shaping.
Future research directions should focus on: (i) extending the model beyond mean-field approximation to incorporate spatial correlations, (ii) developing methods to empirically measure the $R$ and $L$ parameters in real social networks, and (iii) exploring applications in controlling misinformation spread and engineering consensus in complex organizations. The mathematical framework developed here provides a solid foundation for these important next steps in sociophysics research.
A Conclusion
This study presents a significant contribution to the understanding of opinion dynamics within complex systems through a master-node model under an annealing approximation framework. By introducing a chiral formulation of the master node, we were able to uncover a rich phase diagram that includes both continuous and discontinuous transitions between ordered and disordered states. These transitions are highly dependent on initial conditions, particularly in systems with two absorbing states: complete disagreement ($m^{*}=0$) and full consensus ($m^{*}=1$). The identification of tricritical points at $(L,R)=(0,0)$ further highlights the limitations in smoothly connecting disordered and ordered phases, reinforcing the idea that consensus formation is not always a gradual process and can exhibit abrupt shifts under certain conditions.
One of the key theoretical contributions lies in the analytical framework developed, which allows for the characterization of symmetry-breaking effects induced by asymmetric influence. This provides valuable insights into how centralized nodes shape collective outcomes, paving the way for applications in organizational management, public policy, and misinformation control. The stability analysis performed using isocline methods revealed the emergence of a third stationary point $\tilde{m}^{(01)}=\frac{L}{R+L}$, depending on the region of the $(L,R)$ parameter space. This finding enriches our understanding of the role of asymmetry in shaping the final state of the system, acting as a biasing mechanism that can either accelerate or delay consensus depending on the coupling parameters.
We hope this study sheds light on the fundamental mechanisms governing opinion dynamics in structured populations, particularly in systems where central nodes exert asymmetric influence over collective outcomes. By framing opinion formation through the lens of statistical mechanics and phase transitions, we bridge the gap between abstract modeling and practical applications in sociophysics. Looking ahead, future research should aim to extend the current mean-field approximation to include spatial correlations and heterogeneous network structures, bringing the model closer to empirical observations. Moreover, developing methods to measure the $R$ and $L$ parameters in real-world social networks could enable more accurate predictions and interventions in areas such as marketing, political campaigns, and crisis communication.
Ultimately, this line of inquiry contributes to a deeper understanding of how information, influence, and consensus evolve in complex adaptive systems — a question of growing importance in our increasingly interconnected world.
B Appendix – case $\overline{A^{(1)}}=0$ and $A^{(4)}>0$
As mentioned in the main text, the stability analysis of the stationary points are based on the (one-dimensional) isocline method. This means that one investigates the sign of $F_{0}$ in the differential equation $\dot{m}=F_{0}(m)$, which means that the relative positions of the key points 00, $1$, $m_{-}^{(0)}$ and $m_{+}^{(0)}$ are helpful. Although $m=1$ is not a stationary point, it is included in the analysis since the meaningful interval for $m$ is $[0,1]$. The analysis of $m_{-}^{(0)}$ and $m_{+}^{(0)}$ will be shown for reference even if they do not belong to this meaningful interval, though the physical outcome is restricted to the stability of these key points in $[0,1]$ only. The analysis is divided into some parts, as specified in table 3.
Part I Point $(0,0)$
Point $(A^{(4)},-2A^{(4)})$
Point $(0,-A^{(4)})$
Part II Line $R+L+A^{(4)}=0$
Line $L=0$
Curve $R^{2}-4LA^{(4)}=0$
Part III Region (0)
Region (I)
Region (II)
Region (III)
Region (IV)
Region (V)
Region (VI)
Table 3: Division of the $(L,R)$ plane suitable for the construction of the flux diagram of the case $\overline{A^{(1)}}=0$ and $A^{(4)}>0$; the regions in Part III are defined in picture 8.
B.1 Part I (points)
Three situations are investigated in this part.
  • I.1) Point $(0,0)$
    The differential equation for this case is
    $\dot{m}=-A^{(4)}m^{3},$ (B.1)
    and displays a single stationary point at $m^{\ast}=0$.
  • I.2) Point $(A^{(4)},-2A^{(4)})$
    The differential equation for this case is
    $\dot{m}=-A^{(4)}m,$ (B.2)
    and displays a single stationary point at $m^{\ast}=0$.
  • I.3) Point $(0,-A^{(4)})$
    The differential equation for this case is
    $\dot{m}=-A^{(4)}m^{2},$ (B.3)
    and displays a single stationary point at $m^{\ast}=0$.
The three cases above share the same flux diagram, which is depicted in figure 4.
$<$001111
Figure 4: Flux diagram corresponding to the cases I.1, I.2 and I.3.
B.2 Part II (curves)
The stability analysis now are performed on lines and curves.
II.1) Line $R+L+A^{(4)}=0$
The differential equation for this case is
$\dot{m}=\left(L-A^{(4)}\right)m\left(m-\tilde{m}\right),$ (B.4)
where $\tilde{m}:=\frac{L}{L-A^{(4)}}$. The system displays two stationary points at $m^{\ast}=0$ and $m^{\ast}=\tilde{m}$. Since
$\tilde{m}:=\frac{L}{L-A^{(4)}}\left\{\begin{array}[]{llcl}>1&\text{ (meaning $0<1<\tilde{m}$) }&,&L>A^{(4)}\\ \in(0,1)&\text{ (meaning $0<\tilde{m}<1$) }&,&L<0\\ <0&\text{ (meaning $\tilde{m}<0<1$) }&,&0<L<A^{(4)}\end{array}\right.,$ (B.8)
the flux diagram is shown in figure 5.
[…]
Figure 5: Flux diagram - part II.1. The dots represent stationary points and the dashed lines indicate regions outside $[0,1]$.
This case assumes $L\neq A^{(4)}$ and $L\neq 0$; nevertheless, these two situations were analyzed previously (part I.2 and I.3, respectively).
II.2) Line $L=0$
The differential equation for this case is
$\dot{m}=-\left(R+A^{(4)}\right)m^{2}\left(m-\tilde{m}\right),$ (B.9)
where $\tilde{m}:=\frac{R}{R+A^{(4)}}$. The system displays two different stationary points at $m^{\ast}=0$ and $m^{\ast}=\tilde{m}$. Since
$\tilde{m}:=\frac{R}{R+A^{(4)}}\left\{\begin{array}[]{llcl}>1&\text{ (meaning $0<1<\tilde{m}$) }&,&R<-A^{(4)}\\ \in(0,1)&\text{ (meaning $0<\tilde{m}<1$) }&,&R>0\\ <0&\text{ (meaning $\tilde{m}<0<1$) }&,&-A^{(4)}<R<0\end{array}\right.,$ (B.13)
the flux diagram is depicted in figure 6.
[…]
Figure 6: Flux diagram - part II.2. The dots represent stationary points and the dashed lines indicate regions outside $[0,1]$.
This case assumes $R\neq 0$ and $R\neq-A^{(4)}$, but these two scenarios were analyzed previously (parts I.1 and I.3, respectively).
II.3) Curve $R^{2}-4LA^{(4)}=0$
The differential equation for this case is
$\dot{m}=-\frac{\left(R+2A^{(4)}\right)^{2}}{4A^{(4)}}m\left(m-\tilde{m}\right)^{2},$ (B.14)
where $\tilde{m}:=\frac{R}{R+2A^{(4)}}$. The system displays two different stationary points at $m^{\ast}=0$ and $m^{\ast}=\tilde{m}$. Since
$\tilde{m}:=\frac{R}{R+2A^{(4)}}\left\{\begin{array}[]{llcl}>1&\text{ (meaning $0<1<\tilde{m}$) }&,&R<-2A^{(4)}\\ \in(0,1)&\text{ (meaning $0<\tilde{m}<1$) }&,&R>0\\ <0&\text{ (meaning $\tilde{m}<0<1$) }&,&-2A^{(4)}<R<0\end{array}\right.,$ (B.18)
the flux diagram is shown in 7.
[…]
Figure 7: Flux diagram - part II.2. The dots represent stationary points and the dashed lines indicate regions outside $[0,1]$.
This case assumes $R\neq 0$ and $R\neq-2A^{(4)}$ and ; nevertheless, these two situations were analyzed previously (parts I.1 and I.2, respectively).
B.3 Part III (planar regions)
The regions not investigated in Part I and Part II constitute disjoint planar regions, which are displayed in figure 8. In the dark region (region (0) below), there is just a single stationary point, $m^{\ast}=0$, whereas the other six have two more, given by (13). In what follows, the relative positions of the points $\{0,m_{-}^{(0)},m_{+}^{(0)},1\}$ will be examined. These analysis are beneficted by the alternative representation of (13), which is
$m_{\pm}^{(0)}=\frac{2L}{R+2L\mp\sqrt{R^{2}-4LA^{(4)}}}.$ (B.19)
Figure 8: Disjoint regions where the stability analysis is performed separately.
III.0) Region (0): $R^{2}-4LA^{(4)}<0$
This is the dark area in figure 8, and corresponds to the situation where $m_{\pm}^{(0)}$, given in (13), are not real solutions. The differential equation for this case is
$\dot{m}=-m\left[L-\left(R+2L\right)m+\left(R+L+A^{(4)}\right)m^{2}\right],$ (B.20)
where the expression inside the square brackets in (B.20) is strictly positive. As stated previously, and clear from (B.20), there is a single stationary point, which is $m^{\ast}=0$, and the flux diagram is shown in figure 9.
$<$001111
Figure 9: Flux diagram in region (0).
III.1) Region (I): $R^{2}-4LA^{(4)}>0$, $R+2L>\sqrt{R^{2}-4LA^{(4)}}$ and $R>0$
From $0<\sqrt{R^{2}-4LA^{(4)}}<R+2L$, one has $L\left(R+L+A^{(4)}\right)>0$. This implies $L>0$ and $R+L+A^{(4)}>0$, since it is not possible to admit $L<0$ and $R+L+A^{(4)}<0$ at the same time: note that $R+L+A^{(4)}<0$ is equivalent to $(R+2L)+A^{(4)}<L$; the left hand-side is positive ($R+2L>\sqrt{R^{2}-4LA^{(4)}}>0$), while the right-hand side is negative (leading to an impossible situation).
From $R+2L>\sqrt{R^{2}-4LA^{(4)}}$ and $R+L+A^{(4)}>0$, it is straightforward that $0<m_{-}^{(0)}<m_{+}^{(0)}$ from (13). Furthermore, one sees that
$m_{+}^{(0)}=\frac{2L}{2L+\left(R-\sqrt{R^{2}-4LA^{(4)}}\right)}<1.$ (B.21)
Hence, $10<m_{-}^{(0)}<m_{+}^{(0)}<$ 1 in region (I).
III.2) Region (II): $R^{2}-4LA^{(4)}>0$, $R+2L>\sqrt{R^{2}-4LA^{(4)}}$ and $R<0$
From the same reasoning above, one has $R+L+A^{(4)}>0$ and $L>0$, which implies $m_{-}^{(0)}<m_{+}^{(0)}$ by (13). Furthermore, from
$m_{-}^{(0)}=\frac{R+2L-\sqrt{R^{2}-4LA^{(4)}}}{2\left(R+L+A^{(4)}\right)}=\frac{2L}{2L-\left(-R-\sqrt{R^{2}-4LA^{(4)}}\right)},$ (B.22)
it follows that $m_{-}^{(0)}>0$ (from the second member of (B.22)); therefore, it is straightforward that $m_{-}^{(0)}>1$ from the last member of (B.22) (note that $-R-\sqrt{R^{2}-4LA^{(4)}}>0$ but the denominator of the last member is positive because $m_{-}^{(0)}>0$). Hence, $0<1<m_{-}^{(0)}<m_{+}^{(0)}$ in region (II).
III.3) Region (III): $R^{2}-4LA^{(4)}>0$, $R+2L<\sqrt{R^{2}-4LA^{(4)}}$, $R+L+A^{(0)}>0$ and $L<0$
From
$m_{+}^{(0)}=\frac{(-2L)}{(-2L)+\left(\sqrt{R^{2}-4LA^{(4)}}-R\right)},$ (B.23)
since $\sqrt{R^{2}-4LA^{(4)}}-R>0$ for $L<0$, then $10<m_{+}^{(0)}<$ 1. Furthermore,
$m_{-}^{(0)}=\frac{R+2L-\sqrt{R^{2}-4LA^{(4)}}}{2\left(R+L+A^{(4)}\right)}<0,$ (B.24)
since the numerator and denominator do not share the same sign. Hence, $m_{-}^{(0)}<0<m_{+}^{(0)}<1$ in region (III).
III.4) Region (IV): $R^{2}-4LA^{(4)}>0$, $R+2L<\sqrt{R^{2}-4LA^{(4)}}$, $R+L+A^{(0)}>0$ and $L>0$
From $R+L+A^{(0)}>0$, one has $m_{-}^{(0)}<m_{+}^{(0)}$ by (13). Furthermore,
$m_{+}^{(0)}=\frac{2L}{R+2L-\sqrt{R^{2}-4LA^{(4)}}}<0,$ (B.25)
since the numerator and denominator do not share the same sign. Hence, $m_{-}^{(0)}<m_{+}^{(0)}<0<1$ in region (IV).
III.5) Region (V): $R^{2}-4LA^{(4)}>0$, $R+2L<\sqrt{R^{2}-4LA^{(4)}}$, $R+L+A^{(0)}<0$ and $L<0$
From $R+L+A^{(0)}<0$, one has $m_{+}^{(0)}<m_{-}^{(0)}$ by (13). Furthermore,
$m_{+}^{(0)}=\frac{(-2L)}{(-2L)+\left(\sqrt{R^{2}-4LA^{(4)}}-R\right)}\in(0,1),$ (B.26)
since $\sqrt{R^{2}-4LA^{(4)}}-R>0$ for $L<0$. This leads, evidently, to $m_{-}^{(0)}>0$. Finally, from
$m_{-}^{(0)}=\frac{(-2L)}{(-2L)-\left(\sqrt{R^{2}-4LA^{(4)}}+R\right)},$ (B.27)
one sees that the denominator is positive, because of $m_{-}^{(0)}>0$ and the positivity of the numerator. Since $\sqrt{R^{2}-4LA^{(4)}}+R>0$ for $L<0$, it implies $(-2L)>(-2L)-\left(\sqrt{R^{2}-4LA^{(4)}}+R\right)>0$, which leads to $m_{-}^{(0)}>1$. Hence, $0<m_{+}^{(0)}<1<m_{-}^{(0)}$ in region (V).
III.6) Region (VI): $R^{2}-4LA^{(4)}>0$, $R+2L<\sqrt{R^{2}-4LA^{(4)}}$, $R+L+A^{(0)}<0$ and $L>0$
From $R+L+A^{(0)}<0$, one has $m_{+}^{(0)}<m_{-}^{(0)}$ by (13). Furthermore,
$m_{+}^{(0)}=\frac{2L}{R+2L-\sqrt{R^{2}-4LA^{(4)}}}<0,$ (B.28)
since the numerator and denominator do not share the same sign. Moreover, using the representation (B.19), one sees that $m_{-}^{(0)}>0$ because both the numerator and denominator have the same sign. Finally, from
$m_{-}^{(0)}=\frac{2L}{2L-\left(-R-\sqrt{R^{2}-4LA^{(4)}}\right)},$ (B.29)
one sees that the denominator is positive, because of $m_{-}^{(0)}>0$ and the positivity of the numerator. The conditions $R+L+A^{(0)}<0$ and $L>0$ implies $R<0$. Therefore, $-R-\sqrt{R^{2}-4LA^{(4)}}>0$ for $R<0$ and $L>0$, which leads to $02L>2L-\left(-R-\sqrt{R^{2}-4LA^{(4)}}\right)>$ 0, and, consequently, to $m_{-}^{(0)}>1$. Hence, $m_{+}^{(0)}<0<1<m_{-}^{(0)}$ in region (VI).
The flux diagram of Part III (except for region (0), which was already shown) is displayed in figure 10.
[…]
Figure 10: Flux diagram - part III. The dots represent stationary points and the dashed lines indicate regions outside $[0,1]$.
C Appendix - case $\overline{A^{(1)}}>0$ and $A^{(4)}=0$
In this case, the absorbing state $m^{\ast}=1$ is predicted by the model, while the stationary point $m^{\ast}=0$ is excluded by the condition $\overline{A^{(1)}}\neq 0$. In principle, this case would follow the same reasoning as in the previous appendix, but a shortcut is available exploring the symmetry of the model. By introducing the variable
$n(t):=1-m(t),$ (C.1)
which also belongs to the interval $[0,1]$, the dynamical system $\dot{m}=F_{1}(m)$, where $F_{1}(m)$ is given in (25), is cast as
$\dot{n}=-n\left[R-\left(2R+L\right)n+\left(L+R+\overline{A^{(1)}}\right)n^{2}\right],$ (C.2)
where the right-hand side is similar to $F_{0}(m)$ by exchanging $R\leftrightarrow L$ and $A^{(4)}\leftrightarrow\overline{A^{(1)}}$. Furthermore, the roots of the polynomial inside the square brackets in (C.2) are
$n_{\pm}:=\frac{2R+L\pm\sqrt{L^{2}-4R\overline{A^{(1)}}}}{2\left(R+L+\overline{A^{(1)}}\right)},$ (C.3)
which is also similar to $m_{\pm}^{(0)}$ in (13) by exchanging $R\leftrightarrow L$ and $A^{(4)}\leftrightarrow\overline{A^{(1)}}$. It is also important to notice that these two sets of stationary solutions are related by
$m_{\mp}^{(1)}+n_{\pm}=1.$ (C.4)
Hereby, one may benefit from the previous appendix by adapting the results obtained there to the present case, and then express the outcomes in terms of the variables specific to this case. The analysis is again separated into disjoint regions specified in table 4, and the results are listed for reference.
Part IV Point $(0,0)$
Point $(-2\overline{A^{(1)}},\overline{A^{(1)}})$
Point $(-\overline{A^{(1)}},0)$
Part V Line $R+L+\overline{A^{(1)}}=0$
Line $R=0$
Curve $L^{2}-4R\overline{A^{(1)}}=0$
Part VI Region (0)
Region (I)
Region (II)
Region (III)
Region (IV)
Region (V)
Region (VI)
Table 4: Division of the $(L,R)$ plane suitable for the construction of the flux diagram of the case $\overline{A^{(1)}}>0$ and $A^{(4)}=0$; the regions in Part VI are defined in picture 11.
C.1 Part IV (points)
The flux diagram of part IV is given in table 5.
Point $F_{1}(m)$ Flux diagram
$(0,0)$ $\overline{A^{(1)}}\left(1-m\right)^{3}$ 001111$>$
$(-2\overline{A^{(1)}},\overline{A^{(1)}})$ $\overline{A^{(1)}}\left(1-m\right)$
$(-\overline{A^{(1)}},0)$ $\overline{A^{(1)}}\left(1-m\right)^{2}$
Table 5: Flux diagram for the points in part IV.
C.2 Part V (curves)
The flux diagrams for the lines and curves in part V are given in table 6
Curve $F_{1}(m)$ $\tilde{m}^{(1)}$ Condition Flux diagram
[…]
$R>\overline{A^{(1)}}$ $\tilde{m}^{(1)}$001111>>>>>>
$R=0$ $\big{(}L+\overline{A^{(1)}}\big{)}\big{(}1-m\big{)}^{2}\big{(}\tilde{m}^{(1)}-m\big{)}$ $\frac{\overline{A^{(1)}}}{\overline{L+A^{(1)}}}$ $L<-\overline{A^{(1)}}$ $\tilde{m}^{(1)}$001111>>>>>>
$-\overline{A^{(1)}}<L<0$
[…]
$L>0$
[…]
$L^{2}-4R\overline{A^{(1)}}=0$ $\frac{\big{(}2\overline{A^{(1)}}+L\big{)}^{2}}{4\overline{A^{(1)}}}\big{(}1-m\big{)}\big{(}\tilde{m}^{(1)}-m\big{)}^{2}$ $\frac{\overline{2A^{(1)}}}{\overline{2A^{(1)}}+L}$ $L<-2\overline{A^{(1)}}$ $\tilde{m}^{(1)}$001111>>>>>>
[…]
Table 6: Flux diagram for the lines and curve of part V.
C.3 Part VI (planar regions)
The regions indicated in part VI are shown in figure 11, and the flux diagram is shown in table 13.
Figure 11: Disjoint regions where the stability analysis is performed separately.
The dark area in figure 11, and corresponds to the situation where $m_{\pm}^{(1)}$, given in (26), are not real solutions. The differential equation for this case is
$\dot{m}=\left(1-m\right)\left[\overline{A^{(1)}}-\left(2\overline{A^{(1)}}+L\right)m+\left(R+L+\overline{A^{(1)}}\right)m^{2}\right],$ (C.5)
where the expression inside the square brackets in (C.5) is strictly positive. As stated previously, and clear from (C.5), there is a single stationary point, which is $m^{\ast}=1$, and the flux diagram is shown in figure 12.
$>$001111
Figure 12: Flux diagram in region (0).
The remaining regions are analyzed below.
[…]
Figure 13: Flux diagrams for the regions in part VI.
D Appendix - case $\overline{A^{(1)}}=0$ and $A^{(4)}=0$
In this last scenario, where $\overline{A^{(1)}}=0$ and $A^{(4)}=0$, the equation of motion can be cast as $\dot{m}=F_{01}(m)$, where
$F_{01}(m)$ $:=-Lm+\left(R+2L\right)m^{2}-\left(R+L\right)m^{3}$
$=m\left(1-m\right)\left[\left(R+L\right)m-L\right].$ (D.1)
Apart from the stationary points $m^{\ast}=0$ and $m^{\ast}=1$, a third one, $\tilde{m}^{(01)}$, emerges depending on the region $(L,R)$, and is given by
$\tilde{m}^{(01)}:=\frac{L}{R+L}.$ (D.2)
The analysis is separated into the parts shown in table 4
Part VII Line $R=0$
Line $L=0$
Line $R+L=0$
Part VIII Region (I)
Region (II)
Region (III)
Region (IV)
Region (V)
Region (VI)
Table 7: Division of the $(L,R)$ plane suitable for the construction of the flux diagram of the case $\overline{A^{(1)}}=0$ and $A^{(4)}=0$; the regions in Part II are defined in picture 14.
D.1 Part VII (lines)
In Part VII, the stationary points are $m^{\ast}=0$ or $m^{\ast}=1$ only. The flux diagrams of all cases in this part are depicted in table 8.
Line $F_{01}(m)$ Condition Flux diagram
$R=0$ $-Lm\left(1-m\right)^{2}$ $L>0$ 001111$<$
$L<0$ 001111$>$
$L=0$ $Rm^{2}\left(1-m\right)$ $R>0$ 001111$>$
$R<0$ 001111$<$
$R+L=0$ $-Lm\left(1-m\right)$ $L>0$ 001111$<$
$L<0$ 001111$>$
Table 8: Flux diagram for the lines and curve of part V.
D.2 Part VIII (planar regions)
The stability analysis of part VIII consists of analyzing six disjoint regions indicated in figure 14.
Figure 14: Disjoint regions where the stability analysis is performed separately.
In these regions, a third stationary point (apart from $m^{\ast}=0$ and $m^{\ast}=1$), $\tilde{m}^{(01)}$, given by (D.2), emerges. It is straightforward that
$m^{(01)}=\frac{L}{R+L}=\left\{\begin{array}[]{lcl}>1&,&(R+L>0\textnormal{ and }R<0)\textnormal{ or }(R+L<0\textnormal{ and }R>0)\\ \in(0,1)&,&(R+L>0\textnormal{ and }R,L>0)\textnormal{ or }(R+L<0\textnormal{ and }R,L<0)\\ <0&,&(R+L>0\textnormal{ and }L<0)\textnormal{ or }(R+L<0\textnormal{ and }L>0)\end{array}\right.,$ (D.6)
and this characterization of $\tilde{m}^{(01)}$ implies the flux diagrams presented in figure 15.
[…]
Figure 15: Flux diagram - part VIII. The dots represent stationary points and the dashed lines indicate regions outside $[0,1]$.